Translate into algebraic equation
35/5 + 13= 20
1
Expert's answer
2021-08-09T09:14:35-0400

"Translate \space into \space algebraic \space equation\\
solution\\
given\\

\frac{35}{5} \space + \space 13= \space 20...eq(A)\\
we \space want \space to \space change \space it \space into \space algebraic \space equation\\
---------------------------------\\
algebric \space equation \space definition\\
in \space mathematics \space , \space an \space algebraic \space equation \space is \space an \space expression\\
 \space built \space up \space from \space integer \space constants \space , \space \\
variables \space and \space the \space algebraic \space operations.\\
----------------------------------\\
in \space the \space question, \space we \space have \space a \space single \space equation \space \\
then \space we \space can \space create \space only \space single \space variable \space algebraic \space equation\\
----------------------------------\\
case(i) \space if \space we \space take \space 35 \space as \space a \space variable:-\\
substitute \space 35 \space = \space x \space in \space the \space eq(A)\\
we \space get\\
\frac{x}{5}+13=20\\
simplify\\
x+(5×13)=(5×20)\\
x+65=100\\
this \space represents \space an \space algebraic \space equation\\
-------------------------------------\\
case(ii) \space if \space we \space take \space 13 \space as \space a \space variable:-\\
substitute \space 13 \space = \space x \space in \space the \space eq(A)\\
we \space get\\
\frac{35}{5}+x=20\\
simplify\\
35+(5×x)=(5×20)\\
5x+35=100\\
this \space represents \space an \space algebraic \space equation\\
-------------------------------------\\
case(iii) \space if \space we \space take \space 20 \space as \space a \space variable:-\\
substitute \space 20 \space = \space x \space in \space the \space eq(A)\\
we \space get\\
\frac{35}{5}+13=x\\
simplify\\
35+(5×13)=(5×x)\\
35+65=5x\\
100=5x\\
this \space represents \space an \space algebraic \space equation\\
-------------------------------------\\
case(iv) \space if \space we \space take \space 5 \space as \space a \space variable:-\\
substitute \space 5 \space = \space x \space in \space the \space eq(A)\\
we \space get\\
\frac{35}{x}+13=20\\
simplify\\
35+(x×13)=(20×x)\\
35+13x=20x\\

this \space represents \space an \space algebraic \space equation\\

 \space"